OpenStudy (anonymous):

Solve: x^(x^(x^(x^(x^....)))) = 3

ganeshie8 (ganeshie8):

x^3 = 3 u see.. its just this. solve for x now

OpenStudy (shubhamsrg):

we have this simplified form : x^3 = 3 solve for x.. (assuming x tends to infinyt)

OpenStudy (shubhamsrg):

infinity**

OpenStudy (anonymous):

Thats right. But x have not only one solutions.

OpenStudy (shubhamsrg):

how about x^(x^3) = 3 or x^3 (ln x) = ln 3 i dont know how we can solve this one : but this will surely give another value of x

OpenStudy (anonymous):

Not x^(x^3) = 3. x^3 * x^3 =3 Another value of x is 3^(1/9) And actually there are infinite number of solutions of x. out of which x^(1/3) is the largest.

OpenStudy (shubhamsrg):

x^3 . x^3 = 3 ??? you sure ??? i hope its not (x^x).(x^x)..... ?

OpenStudy (anonymous):

Sorry. My mistake

OpenStudy (shubhamsrg):

hmm,,so that requires the knowledge of interpolation to solve for further values of x. and yes,,must be infinite values of x

OpenStudy (anonymous):

x^(3^3)

OpenStudy (anonymous):

x^(3^(3^(3)))

OpenStudy (anonymous):

and so on

OpenStudy (shubhamsrg):

if you have already substitued that much part to be 3..you have nothing left(ofcorse) i mean 3^3 is not valid..and also further ones are not valid

ganeshie8 (ganeshie8):

u asking if this can look like this : \[\huge{x^{x^3} = 3 }\] ?

OpenStudy (anonymous):

infinity never ends

OpenStudy (shubhamsrg):

your logic is wrong buddy @sauravshakya ,,please re-think.. zabardasti solve kar rahe ho..

OpenStudy (anonymous):

3^(1/9) and 3^(1/27) both are valid

OpenStudy (shubhamsrg):

ofcorse not!! hmm,,

ganeshie8 (ganeshie8):

i think she is correct.. @shubhamsrg .. can you explain my above expression... im confused after putting that :(

OpenStudy (shubhamsrg):

x^(x^3) = 3..yes thats fine.. and not x^(3^3) <---this is humbug! :P

ganeshie8 (ganeshie8):

i mean, her question is valid... 1/9 is ofcourse not equal to 1/27 !!

ganeshie8 (ganeshie8):

ahh.. i see now... thanks @shubhamsrg ;)

OpenStudy (anonymous):

I will return..... With a proof

OpenStudy (shubhamsrg):

hmmm..

OpenStudy (anonymous):

The proof is simple. Since x^(x^(x^(x^(x^....)))) = 3, s^[x^(x^(x^(x...))))] =3, or x^3 =3